# Quantum optical neural networks with programmable nonlinearities

### v1

@@ Abstract
Parametrized quantum circuits are essential components of variational quantum algorithms. Until now, optical implementations of these circuits have relied solely on adjustable linear optical units. In this study, we demonstrate that using programmable nonlinearities, rather than linear optics, offers a more efficient method for constructing quantum optical circuits — especially quantum neural networks. This approach significantly reduces the number of adjustable parameters needed to achieve high-fidelity operation. Specifically, we explored a quantum optical neural network (QONN) architecture composed of meshes of two-mode interferometers programmable by adjustable Kerr-like nonlinearities. We assessed the capabilities of our quantum optical neural network architecture and compared its performance to previously studied architectures that use multimode linear optics units. Additionally, we suggest future research directions for improving programmable quantum optical circuits.
@@ Introduction
Our work draws inspiration from significant experimental progress made in developing nonlinear elements that are sensitive to single-photon-level quantum states [18, 19, 20, 21, 22]. These elements can effectively introduce photon-photon interactions, which are lacking in linear optics [23]. Additionally, similar approaches are being explored in the microwave domain using mature superconducting components, enjoying accessible strong nonlinearities required for the single-photon-sensitive elements, in particular, deterministic nonlinear gates [24]. These advances are motivating the exploration of quantum computing architectures that go beyond traditional linear-optical methods.
Quantum optical neural networks (QONNs) are a specific class of photonic circuits that alternate between multimode linear-optical evolution and single-site nonlinear elements, in order to manipulate photonic quantum states. Similar to classical neural networks, where training involves adjusting linear layers [25], previous research on QONNs has focused on training through adjustable multimode linear optics [26, 27]. In this paper, we introduce a novel approach to parameterized quantum optical circuits, specifically QONNs, which relies on programming single-site nonlinearities instead of programmable multimode linear optics. In particular, we study quantum optical circuits composed of nonlinear Mach-Zehnder interferometer (NMZI) meshes, where the only control parameters are the strengths of single-mode Kerr-like nonlinearities. We assess these circuits on a range of tasks, including generating entangled qubit states, performing quantum logic operations, and implementing variational quantum algorithms.
Our results show that QONNs based on this approach require significantly fewer adjustable parameters to achieve high-quality performance compared to those trained using linear-optics. Moreover, for certain tasks, NMZI-based circuits not only reduce the number of parameters but may also lead to shallower circuit depths. Since single-photon-sensitive nonlinearities are expected to be the primary source of noise, this reduction translates into improved fidelities of QONNs. Thus, our findings suggest that variational quantum optical circuits utilizing programmable single-photon-sensitive nonlinearities offer substantial advantages.
@@ I.1 Variational core of the QONN
Without NMZI blocks introduced in our QONN, the class of possible transformations is limited to the one of multimode linear optical networks — a very narrow class of operations. However, NMZI transformations result in states that are incompatible with single- and dual-rail qubit encodings which are standard methods of photonic information encoding. Neither, the NMZI cannot perform single-qubit operations due to non-adjustable $\varphi_{b}$. For example, Fig. 2 illustrates the four-mode optical circuit utilizing an NMZI, the bulk part of our NMZI mesh (Fig. 1b) consists of. At specific parameter values, $\chi_{1}=\chi_{2}=\pi$ and $\varphi_{b}=0$, this is simply the CZ-gate [15] acting on dual-rail-encoded states at the circuit inputs. However, other choices of $\chi_{1}$, $\chi_{2}$, and $\varphi_{b}$ lead to deviations from the dual-rail qubit encoding. In this work, we show that multiple NMZIs constituent the nonlinear variational core can perform qubit operations.
@@ I.2 Linear optical single-qubit corrections
$\begin{split}U_{\text{MZI}}(\theta_{1},\theta_{2})=U_{\text{DC}}\left(\begin{array}[]{cc}e^{i\theta_{2}}&0\\ 0&1\end{array}\right)U_{\text{DC}}\times\\ \left(\begin{array}[]{cc}e^{i\theta_{1}}&0\\ 0&1\end{array}\right)=\left(\begin{array}[]{cc}e^{i\theta_{1}}\cos(\frac{\theta_{2}}{2})&i\sin(\frac{\theta_{2}}{2})\\ ie^{i\theta_{1}}i\sin(\frac{\theta_{2}}{2})&\cos(\frac{\theta_{2}}{2})\end{array}\right)\end{split}$ (4)
@@ I.4 Relevant parameters of QONNs
In addition to the high-quality realization of a target functionality by the QONNs, we focus closely on two factors: (1) the circuit depth $D$ and (2) the number of adjustable parameters $P$ required to achieve high-quality performance. By circuit depth $D$, we refer to the number of nonlinear layers $L$ in the programmable circuit, based on the reasonable assumption that nonlinearities are the primary contributors to noise and errors. Therefore, it is generally preferable to select circuits with the fewest possible nonlinear layers.
The number of parameters in a quantum neural network can be linked to the computational resources needed for training, as increasing the parameter count may lead to more difficult training due to the curse of dimensionality, which is responsible for the barren plateau phenomenon [29, 30]. However, we recognize that the trainability issue is more complex, and the difficulty of training parameterized quantum circuits largely depends on their underlying architecture. A detailed investigation into the trainability of QONNs will be conducted in a future study.
@@
It is important to note that the scaling of the number of adjustable parameters $P$ with the circuit depth $D$ differs significantly between the proposed nonlinearity-adjustable QONNs and traditional linear-optics-adjustable QONNs. Specifically, adding one nonlinear layer to the former increases the number of adjustable parameters by $M$, whereas in the latter it increases by $M^{2}$. Consequently, the total number of adjustable parameters for the nonlinearity-adjustable QONN is $P_{\text{NL}}=D\times(M-2)+2\lfloor D/2\rfloor$ (excluding the phase shifts used for single-qubit corrections), while for the linear-optics-adjustable QONN, it is $P_{\text{L}}=(D+1)\times M(M-1)$.
@@ II.1 Generation of entangled qubit states
For all the problems considered in this work, the networks were trained using a numerical optimization procedure detailed in Appendix B. During training, the optimization algorithm searched for a global minimum of a cost function (CF) that quantifies the error in the task under consideration. Specifically, for the state generation problem, the cost function is the infidelity: $CF=1-F_{\alpha}(\bm{\chi},\bm{\theta})$. The same optimization procedure was used for QONNs programmed with linear optics and static nonlinearities (see Fig. 3), allowing us to compare these two approaches to constructing QONNs. We analyzed the effect of depth $D$ on the QONN’s capabilities by training the QONN across a range of $D$. Typically, the likelihood of CF converging to a good minimum in a single optimization run is quite low due to the algorithm often falling into a local minimum, which necessitated multiple optimization runs for each target state $|G_{n}(\alpha)\rangle$ and depth value $D$.
@@
Figure 4: Preparation of the $3$-qubit states (II.1) by different QONNs: state infidelity $1-F$ for the QONN with programmable nonlinearities as a function of circuit depth $D$ for various values of $\alpha$ and different biases $\varphi_{b}=0$ (a), $\pi/4$ (b), $\pi/2$ (c), $43\pi/$ 4 (d) and $\pi$ (e). The same dependence for QONN with programmable linear optics is plotted in (f).
Figure 5: Preparation of the $4$-qubit states (II.1) by different QONNs. Parameters are the same as in Fig. 4.
Fig. 4 and Fig. 5 show the obtained dependencies of infidelity on circuit depth for various values of the state parameter $\alpha$ for $n=3$ and $n=4$ qubits, respectively. In addition to varying $\alpha$, we also examined different phase shifts $\varphi_{b}$ that control the linear optical part of the QONN transformation. We selected linear optical configurations with $\varphi_{b}$ values of 0, $\pi/4$, $\pi/2$, $43\pi/$ 4, and $\pi$ (the corresponding dependencies are drawn in plots a)-e) of the figures). Our QONN is not capable of preparing the states for all presented values of $\varphi_{b}$ — only two of these values, specifically $\varphi_{b}=\pi/4$ and $\varphi_{b}=\pi$, work. We notice that the dependencies are radically different for these two cases. At $\varphi_{b}=\pi/4$, the infidelity almost monotonically decreases with increasing circuit depth for all target states, so one predefined depth value $D$ is universal for all the states. In contrast, at $\varphi_{b}=\pi$ infidelity for the $3$-qubit states is non-monotonic and also each target state requires its own optimal value of $D$. At the same time, for the $4$-qubit states infidelity drops monotonically and starting from a value of $D=10$ all the states can be prepared with high quality.
@@
Figure 6: Preparation of parametrized entangled $3$-qubit states (II.1) without single-qubit corrections: the state infidelity $1-F$ as a function of circuit depth $D$ for entangled states preparation with various parameter $\alpha$ at different phase-shift biases $\varphi_{b}=0$ (a) and $\pi$ (b).
@@
The infidelities for the QONN programmed by linear optics are also plotted in Fig. 4f and Fig. 5f in order to compare their performance with our QONNs. As one can see, the QONN programmed by linear optics is capable of preparing the states as well. Moreover, it requires lower number of (static) nonlinear layers than in our QONNs. Importantly, however, that the number of adjustable phase-shifts $P_{L}$ required to prepare the states with high fidelity scales quadratically, $P_{L}\propto M^{2}$, with the number of modes $M=2n$, rather than linearly, $P_{NL}\propto M$, as in our QONN. Therefore, $P_{NL}\ll P_{L}$ for large qubit numbers $n$. In particular, for maximally entangled states of $n=3,4$ qubits the parameter counts reads $P_{NL}=4,62$ and $P_{L}=60,168$ for our QONN and the counterpart programmed by linear optics, respectively (see Table 1).
@@
Table 1: Circuit depth $D$ and parameter count $P$ of QONNs peforming maximally entangled state generation (see Fig. 45). NL and LO are the QONNs with adjustable nonlinearities and the counterparts trained by linear optics, respectively.
@@
cell11 = r=2, vlines, hline1,3,6 = -, hline2 = 2-5, & $D$ $P$
$3$-/$4$-qubit $3$-/$4$-qubit
NL ($\varphi_{b}=0$, $\alpha=\pi/4$) 1/– 4/–
NL ($\varphi_{b}=\pi$, $\alpha=\pi/4$) 5/9 24/62
LO 1/2 60/168
Figure 7: Preparation of randomly sampled $3$-qubit states: the state infidelity $1-F$ as a function of circuit depth $D$ at different values of linear phase-shift biases $\varphi_{b}$ for the proposed QONN (solid curves) and for the QONN programmed by linear optics (dashed curve). The curves have been obtained for $30$ randomly generated states.
@@
It is instructive to compare the number of adjustable parameters required by the QONNs relative to the values set by circuits capable of preparing arbitrary qubit states. One can use the following obvious reasoning: a generic $n$-qubit state can be parameterized by $12^{n+1}-$ 1 real parameters (if common phase is not included), suggesting that a parameterized quantum circuit designed to prepare arbitrary states should include at least that much programmable parameters. In particular, for $n=3$ the universal circuit should contain $P\geq{}15$ parameters, respectively. At the same time, using the results depicted in Fig. 7, one obtains $P_{NL}=54$ and $P_{L}=150$ for our and the linear-optically programmed QONN, respectively. This suggests that both types of parameterized quantum optical circuits are far suboptimal from the gate-based circuits. However, using programmable nonlinearities rather than programmable linear optics turns out to be beneficial for the state generation task.
Table 2: Parameters of QONNs peforming $3$-qubit state generation (see Fig. 7). NL and LO are the QONNs with adjustable nonlinearities and the counterparts trained by linear optics, respectively.
@@
vlines, hline1-2 = -, hline4 = -, & $D$ $P$
NL ($\varphi_{b}=\pi$) 11 54
LO 4 150

### v2

@@ Abstract
Parametrized quantum circuits are essential components of variational quantum algorithms. Until now, optical implementations of these circuits have relied solely on adjustable linear optical units. In this study, we demonstrate that using programmable nonlinearities, rather than linear optics, offers a more efficient method for constructing quantum optical circuits — especially quantum neural networks. This approach significantly reduces the number of adjustable parameters and the circuit depth needed to achieve high-fidelity operation. Specifically, we explored a quantum optical neural network (QONN) architecture composed of meshes of two-mode interferometers programmable by adjustable Kerr-like nonlinearities. We assessed the capabilities of our quantum optical neural network architecture and compared its performance to previously studied architectures that use multimode linear optics units. Additionally, we suggest future research directions for improving programmable quantum optical circuits.
@@ Introduction
Our work draws inspiration from significant experimental progress made in developing non-linear elements that are sensitive to single-photon-level quantum states [18, 19, 20, 21, 22, 23], especially those compatible with integrated photonic circuits [20, 21, 22, 23]. These elements can effectively introduce photon-photon interactions, which are lacking in linear optics [24]. Additionally, similar approaches are being explored in the microwave domain using mature superconducting components, enjoying accessible strong nonlinearities required for the single-photon-level elements, in particular, deterministic nonlinear gates [25]. These advances are motivating the exploration of quantum computing architectures that go beyond traditional linear-optical methods.
Quantum optical neural networks (QONNs) are a specific class of photonic circuits that alternate between multimode linear-optical evolution and single-site nonlinear elements, in order to manipulate photonic quantum states. Similar to classical neural networks, where training involves adjusting linear layers [26], previous research on QONNs has focused on training through adjustable multimode linear optics [27, 28]. In this paper, we introduce a novel approach to parameterized quantum optical circuits, specifically QONNs, which relies on programming single-site nonlinearities instead of programmable multimode linear optics. In particular, we study quantum optical circuits composed of nonlinear Mach-Zehnder interferometer (NMZI) meshes, where the only control parameters are the strengths of single-mode Kerr-like nonlinearities. We assess these circuits on a range of tasks, including generating entangled qubit states, performing Bell-state discrimination, and implementing variational quantum algorithms.
Our results show that QONNs based on this approach require significantly fewer adjustable parameters and programmable optical element layers to achieve high-quality performance compared to those trained using linear-optics. This reduction translates into improved fidelities of QONNs. Thus, our findings suggest that variational quantum optical circuits utilizing programmable single-photon-sensitive nonlinearities offer substantial advantages.
@@ I.1 Variational core of the QONN
Without nonlinearities in the NMZI blocks the class of possible transformations is limited to the one of multimode linear optical networks, which is a very narrow class of operations in the photonic Hilbert space. Generally, NMZI transformations operating in the full Hilbert space result in states that are incompatible with single- and dual-rail qubit encodings which are standard methods of photonic information encoding. Also, the NMZI cannot perform single-qubit operations due to non-adjustable $\varphi_{b}$. For example, Fig. 2 illustrates the four-mode optical circuit utilizing an NMZI. At specific parameter values, $\chi_{1}=\chi_{2}=\pi$ and $\varphi_{b}=0$, this is simply the CZ-gate [15] acting on dual-rail-encoded states at the circuit inputs. However, other choices of $\chi_{1}$, $\chi_{2}$, and $\varphi_{b}$ lead to deviations from the dual-rail qubit encoding. In this work, we show that multiple NMZIs constituent the nonlinear variational core can perform qubit operations.
@@ I.2 Linear optical single-qubit corrections
$\begin{split}U_{\text{MZI}}(\theta_{1},\theta_{2})=U_{\text{DC}}\left(\begin{array}[]{cc}e^{i\theta_{2}}&0\\ 0&1\end{array}\right)U_{\text{DC}}\left(\begin{array}[]{cc}e^{i\theta_{1}}&0\\ 0&1\end{array}\right)=\\ =e^{i\theta_{2}/2}\left(\begin{array}[]{cc}e^{i\theta_{1}}\cos(\theta_{2}/2)&i\sin(\theta_{2}/2)\\ ie^{i\theta_{1}}\sin(\theta_{2}/2)&\cos(\theta_{2}/2)\end{array}\right)\end{split}$ (4)
@@ I.4 Relevant parameters of QONNs
Currently, integrated photonics technologies represent a promising avenue for the development of quantum optical circuits [30, 31]. Several key components of these circuits have already been demonstrated on-chip, including photon sources [32, 33], photodetectors [34], programmable linear optics [12], and single-photon-sensitive nonlinearities [20, 21]. However, only a subset of these elements have been integrated into a single platform, and much work remains to fully integrate all the necessary components for large-scale quantum optical circuits.
Our work advances the scalability of programmable quantum optical circuits, which typically consist of two main components: multimode linear optics, commonly implemented using MZI meshes, and single-photon-sensitive nonlinearities. In practical implementations, each type of the components introduces different degrees of losses, errors, and noise, and each poses different fabrication challenges. While it is generally desirable to minimize the number of nonlinear layers, analyzing the impact of the nonidealities introduced by these components in a QONN is complex and requires specific details about the underlying platform.
Linear MZI meshes are relatively simple to fabricate, as they consist of standard elements such as static directional couplers and programmable phase-shifters [8]. However, constructing even moderately sized universal programmable linear-optical circuits remains challenging due to their large size and relatively slow modulation speeds [12, 35]. Recent proposals suggest overcoming these limitations by replacing traditional phase-shifting elements with quantum dots as reconfigurable phase shifters [36]. This approach aims to eliminate the need for bulky linear phase shifters, albeit at the cost of more complex fabrication and control.
Single-photon-sensitive nonlinearities can be realized using single emitters, such as quantum dots, embedded in photonic waveguides or cavities. The strong confinement of light in these structures enables near-unity light-matter coupling efficiencies. Recently, a nonlinear phase shift was demonstrated using a single quantum dot within a planar nanophotonic waveguide [21]. It is worth noting that making nonlinear elements programmable is no more difficult than realizing static nonlinear elements. Since the variable parameter of a nonlinearity is determined by the coupling strength, it can be adjusted post-fabrication, for example, by detuning the emitter’s transition frequency or the resonant frequency of the cavity.
To highlight the advantages of our QONN architecture, we focus on two key characteristics of quantum optical circuits: (1) the circuit depth $D$ and (2) the number of adjustable parameters $P$ required to achieve high-quality performance. Specifically, we examine the linear-optical circuit depth $D_{\text{LO}}$, defined as the number of linear phase-shift layers in the variational core, and the nonlinear circuit depth $D_{\text{NL}}$, which corresponds to the number of nonlinear layers. It is important to note that the proposed nonlinearity-adjustable QONNs have $D_{\text{LO}}=0$, as they do not have programmable linear phase shifts in their variational core. In contrast, traditional linear-optics-adjustable QONNs have both $D_{\text{LO}}$ and $D_{\text{NL}}$ greater than zero, and these two parameters are interrelated. For instance, assuming a compact version of the universal MZI meshes with the minimum number of phase-shift layers [37], the relationship between $D_{\text{LO}}$ and $D_{\text{NL}}$ and is given by:
$D_{\text{LO}}=(M+1)(D_{\text{NL}}+1).$ (5)
In the following, we use $D_{\text{NL}}$ to quantify the circuit depths for both QONN architectures and additionally compute $D_{\text{LO}}$ for the linear-optics-adjustable QONNs using the above equation.
@@
The number of parameters $P$ in a quantum neural network is related to the computational resources required for training. Increasing the number of parameters can make training more challenging due to the curse of dimensionality, which is associated with the barren plateau phenomenon [38, 39]. However, the difficulty of training is more nuanced, and it largely depends on the architecture of the quantum circuit. A detailed analysis of the trainability of QONNs will be conducted in future work. The scaling of the number of adjustable parameters $P_{\text{NL}}/P_{\text{LO}}$ with circuit depth $D_{\text{NL}}$ differs significantly between the proposed nonlinearity-adjustable QONNs and traditional linear-optics-adjustable QONNs. Specifically, adding one nonlinear layer to the former increases the number of adjustable parameters by MM, while in the latter, the increase is proportional to $M(M-1)$. As a result, the total number of adjustable parameters for the nonlinearity-adjustable QONN is given by $P_{\text{NL}}=D_{\text{NL}}\times(M-2)+2\lfloor D_{\text{NL}}/2\rfloor$ excluding the phase shifts used for single-qubit corrections. For the linear-optics-adjustable QONN, the total number of parameters is $P_{\text{L}}=(D_{\text{NL}}+1)\times M(M-1)$.
@@ II.1 Generation of entangled qubit states
For all the problems considered in this work, the networks were trained using a numerical optimization procedure detailed in Appendix B. During training, the optimization algorithm searched for a global minimum of a cost function (CF) that quantifies the error in the task under consideration. Specifically, for the state generation problem, the cost function is the infidelity: $CF=1-F_{\alpha}(\bm{\chi},\bm{\theta})$. The same optimization procedure was used for QONNs programmed with linear optics and static nonlinearities (see Fig. 3), allowing us to compare these two approaches to constructing QONNs. We analyzed the effect of depth $D$ on the QONN’s capabilities by training the QONN across a range of $D$. Typically, the likelihood of CF converging to a good minimum in a single optimization run is quite low due to the algorithm often falling into a local minimum, which necessitated multiple optimization runs for each target state $|G_{N}(\alpha)\rangle$ and depth value $D_{\text{NL}}$.
@@
Figure 4: Preparation of the $3$-qubit states $|\text{G}_{N}(\alpha)\rangle$ ($N=3$) with various entanglement degree by different QONNs: state infidelity $1-F$ for the QONN with programmable nonlinearities as a function of circuit depth $D_{\text{NL}}$ for various values of $\alpha$ and different biases $\varphi_{b}=0$ (a), $\pi/4$ (b), $\pi/2$ (c), $43\pi/$ 4 (d) and $\pi$ (e). The same dependence for QONN with programmable linear optics is plotted in (f).
@@
Figure 5: Preparation of the $4$-qubit states $|\text{G}_{N}(\alpha)\rangle$ ($N=4$) with various entanglement degree by different QONNs: state infidelity $1-F$ for the QONN with programmable nonlinearities as a function of circuit depth $D_{\text{NL}}$ for various values of $\alpha$ and different biases $\varphi_{b}=0$ (a), $\pi/4$ (b), $\pi/2$ (c), $43\pi/$ 4 (d) and $\pi$ (e). The same dependence for QONN with programmable linear optics is plotted in (f).
Fig. 4 and Fig. 5 show the obtained dependencies of infidelity on circuit depth for various values of the state parameter $\alpha$ for $N=3$ and $N=4$ qubits, respectively. In addition to varying $\alpha$, we also examined different phase shifts $\varphi_{b}$ that control the linear optical part of the QONN transformation. We selected linear optical configurations with $\varphi_{b}$ values of 0, $\pi/4$, $\pi/2$, $43\pi/$ 4, and $\pi$ (the corresponding dependencies are drawn in plots a)-e) of the figures). Our QONN is not capable of preparing the states for all presented values of $\varphi_{b}$ — only two of these values, specifically $\varphi_{b}=\pi/4$ and $\varphi_{b}=\pi$, work. We notice that the dependencies are radically different for these two cases. At $\varphi_{b}=\pi/4$, the infidelity almost monotonically decreases with increasing circuit depth for all target states, so one predefined depth value $D_{\text{NL}}$ is universal for all the states. In contrast, at $\varphi_{b}=\pi$ infidelity for the $3$-qubit states is non-monotonic and also each target state requires its own optimal value of $D_{\text{NL}}$. At the same time, for the $4$-qubit states infidelity drops monotonically and starting from a value of $D_{\text{NL}}=10$ all the states can be prepared with high quality.
Figure 6: Preparation of the $3$-qubit states $|\text{G}_{N}(\alpha)\rangle$ ($N=3$) with various entanglement degree by the linear-optically programmed QONN without single-qubit corrections: the state infidelity $1-F$ as a function of circuit depth $D_{\text{NL}}$ for entangled states preparation with various parameter $\alpha$ at different phase-shift biases $\varphi_{b}=0$ (a) and $\pi$ (b).
@@
The infidelities for the QONN programmed by linear optics are also plotted in Fig. 4f and Fig. 5f in order to compare their performance with our QONNs. As one can see, the QONN programmed by linear optics is capable of preparing the states as well. Moreover, it requires a lower number of (static) nonlinear layers $D_{\text{NL}}$ than in our QONNs. Importantly, however, that the number of phase-shift layers $D_{\text{LO}}$ and the total number of adjustable phase-shifts $P_{\text{LO}}$ required to prepare the states with high fidelity scales is much larger, than the number of nonlinear phase-shift layers $D_{\text{NL}}$ and the total number of adjustable nonlinear phase-shifts $P_{\text{NL}}$ in our QONN.
In particular, for maximally entangled states of $N=3,4$ qubits the parameter counts reads $P_{\text{NL}}=4,62$ and $P_{\text{LO}}=60,168$ for our QONN and the counterpart programmed by linear optics, respectively. Table II.1 summarizes the obtained QONN parameters. The much smaller nonlinear depth $D_{\text{NL}}$ of our QONN than the linear depth $D_{\text{LO}}$ of the counterpart QONN suggests that it can be implemented on a compact quantum circuit.
@@
Table 1: Nonlinear $D_{\text{NL}}$ and linear circuit depth $D_{\text{LO}}$ and parameter count $P$ for QONNs peforming maximally entangled state generation (see Fig. 45). NL and LO are the QONNs with adjustable nonlinearities and the counterparts trained by linear optics, respectively.
@@
cell11 = r=2, vlines, $D_{\text{NL}}$ $D_{\text{LO}}$ $P$
$3$-/$4$-qubit $3$-/$4$-qubit $3$-/$4$-qubit
NL ($\varphi_{b}=0$) 1/– 0 4/–
NL ($\varphi_{b}=\pi$) 5/9 0 24/62
LO 1/2 14/27 60/168
Figure 7: Preparation of randomly sampled $3$-qubit states: the state infidelity $1-F$ as a function of circuit depth $D_{\text{NL}}$ at different values of linear phase-shift biases $\varphi_{b}$ for the proposed QONN (solid curves) and for the QONN programmed by linear optics (dashed curve). The curves have been obtained for $30$ randomly generated states.
@@
It is instructive to compare the number of adjustable parameters required by the QONNs relative to the values set by circuits capable of preparing arbitrary qubit states. One can use the following obvious reasoning: a generic $N$-qubit state can be parameterized by $12^{N+1}-$ 1 real parameters (if common phase is not included), suggesting that a parameterized quantum circuit designed to prepare arbitrary states should include at least that much programmable parameters. In particular, for $N=3$ the universal circuit should contain $P\geq{}15$ parameters, respectively. At the same time, using the results depicted in Fig. 7, one obtains $P_{\text{NL}}=54$ and $P_{\text{LO}}=150$ for our and the linear-optically programmed QONN, respectively. This suggests that both types of parameterized quantum optical circuits are far suboptimal from the gate-based circuits. However, using programmable nonlinearities rather than programmable linear optics turns out to be beneficial for the state generation task.
Table II.1 summarizes the QONN circuit parameters obtained. As before, the depth of our QONN with adjustable nonlinearities turned out to be smaller, than the depth of the QONN programmed by linear optics.
Table 2: Nonlinear $D_{\text{NL}}$ and linear circuit depth $D_{\text{LO}}$ and parameter count $P$ of QONNs peforming $3$-qubit state generation (see Fig. 7). NL and LO are the QONNs with adjustable nonlinearities and the counterparts trained by linear optics, respectively.
@@
vlines, $D_{\text{NL}}$ $D_{\text{LO}}$ $P$
NL ($\varphi_{b}=\pi$) 11 0 54
LO 4 35 150